How many grams of H₂ can be formed from 89.0 grams of NH₃ in the following reaction?

2 NH₃(g) → 3 H₂ (g) + N₂ (g)


1
Expert's answer
2021-10-05T01:44:19-0400

Molar mass of NH3= 17.031

89/17.031= 5.23moles

Mole ratio of NH3:H2

2:3

= 5.23×3/2

= 7.84mol

Molar mass of H2 = 2.016 g/mol

2.016×7.84= 15.8g